### v1

# Through the Looking Glass: Bitcoin Treasury Companies

Abstract
‘Bitcoin treasury companies’ have taken stock markets by storm amassing billions of dollars worth of tokens. More than a hundred companies active in various crypto-tokens, not just Bitcoin, are listed. The ability to generate leverage, for example stemming from borrowing collateralised by stock portfolios, helps to understand this phenomenon.
In addition, the extension of the binary-choice Kelly criterion to incorporate uncertainty in the form of the Kullback-Leibler divergence or more generally Bregman divergence is studied.
I Introduction
The Bitcoin whitepaperbtc dates back to October 2008, the height of the ‘great financial crisis’ (GFC) and only about a month after the Lehman bankruptcy. Legacy finance, as it is now sometimes called, was in disarray and a search for alternatives gathered pace. Bitcoin, maybe initially more of a Gedankenexperiment than a practical means of exchange, has over the intervening years grown dramatically. The associated coin BTC now has a market capitalisation of more than two trillion. Thousands of other coins have joined the proverbial gold rush. For years legacy finance largely ignored these upstarts, but the growth of BTC’s and other tokens’ market capitalisations has piqued curiosity and greed.
Companies, and not just individuals, have started hoarding bitcoins. There are currently over a hundred companies that are devoted to this one task - the holding of crypto-assets. Many market themselves as vehicles whose mission it is to acquire ever-more and ever-faster crypto-tokens financed by a mixture of equity and various form of debt issuance. Puzzlingly, their market capitalisation can be significantly higher than the net asset value of their crypto-assets. This is especially surprising when the non-crypto business of these companies is often between insignificant and non-existent. How can this be? The paper suggests a possible answer, which goes beyond irrational exuberance.
In the next section popular new performance indicators are briefly introduced and in the subsequent section the role of leverage is explored. The penultimate section of the paper is devoted to an extension of the one-dimensional Kelly criterion to cases where the winning probability is uncertain. A conclusion rounds of the paper.
II Key performance indicators ‘reimagined’
One of the key performance indicators (KPIs) touted by crypto-treasury companies is mNAV, which is the ratio of enterprise value [^1] In the case of an ‘idealised’ company, which is the only company we consider, the enterprise value is equivalent to the market capitalisation. We assume the company is debt free, no preferred equity has been issued and (for completeness sake) no pension liabilities exist. to the value of the crypto-tokens held. If this is above one issuing equity to purchase further tokens is value accretive. To see this take a simple example. Assume the token trades at \$$\$1$, and the mNAV is $4$. Further assume $4$ shares have been issued trading at \$$\$1$ each, and one token is in the companies treasury. Next a share is issued at the prevailing price of \$$\$1$ and one token is bought. Assuming an unchanged mNAV, the price per share becomes \$ $\$\,8/5$ an increase of $80\%$. Magically, everyone seems to be better off. How can this be? Is it due to a suspension of disbelief, since the gap between net asset value (NAV) and the market capitalisation has widened [^2] There are also tax and other hurdles that prevent investors from buying tokens directly. For example, in Japan holding Bitcoin can be less tax efficient than stock ownership. Cryptocurrency capital gains tax in Japan is up to $55\%$, while the comparable tax on stocks is only about $20\%$. These tax rates are currently under review.. Why overpay for fractional ownership of crypto-tokens via an intermediary, when either through direct ownership or some other intermediary, as should be possible in a competitive market, the same exposure can be obtained cheaper?
‘Bitcoin per share’, the fraction of Bitcoin held per share, is also a popular KPI. A rise ‘Bitcoin per share’ is viewed as beneficial for the shareholders. This can be achieved with debt or with equity issuance, as we saw above, for mNAV larger than one. Debt issuance is interesting, since it enables the shift of risk to outside parties. This is also true for borrowing by shareholders against portfolio value. The paper considers in the next section credit transfers by borrowing of shareholders but not by the company itself. The different types of credit transfers are not interchangeable, even if in both cases risk is shifted away from the ‘centre’.
Other KPIs including BTC yield, BTC gain and BTC \$ gain have also been bandied about but are maybe not as helpful, or at least add little beyond the metrics mentioned above, to understand the price behaviour. Once mNAV approaches one, other KPIs as well as price impact [^3] For an analysis of strategies incorporating price impact see bkm2016 . Here we are interested with the price impact on Bitcoin by companies buying a substantial amount of the outstanding tokens. gain in importance. In the next section we return to mNAV.
III Mirror images: Is it a virtuous, a vicious or a virtual cycle?
In section II of a paper on ‘rational market failures’ meister2022 a cyclical strategy was described that allowed risk to be passed from stock investors to credit providers. With sufficient market impact an investor was able to extract money. Next, an outline of how the process works. The investor would buy a particular stock, drive up the stock price, since one-sided demand can lift the price, and the portfolio value. These portfolio gains the investor could use as collateral to borrow against. This would lead to a virtuous cycle for the stockholders, and a potentially vicious cycle for the lenders, which would accumulate risk. In extreme cases the investor could extract cash and come out ahead, even if the stock eventually crashes. For details, see the earlier paper.
The ‘innovation’ of crypto-treasury companies is to exploit the counter-intuitive situation of having a mNAV larger than one. In its simplest form the issuance of stock converted back into crypto-tokens raises the stock price, followed by more stock issuance. Additional stock in one scenario is bought by existing holders of the stock, who borrow against the ever increasing value of their stock portfolio to finance the new purchases. This is true even without any overt price impact on crypto-tokens. It wouldn’t be surprising if the steady buying by treasury companies of an assets that is in the case of Bitcoin of fixed supply will in addition raise the price of the token. The efficiency and thereby viability of the equity-issuance strategy only depends on banks or other credit providers treating stock of Bitcoin treasury companies like other assets that can be deposited as collateral for loans with a reasonable haircut. When collateral requirements are raised, the attractiveness of these companies as endlessly equity-issuing financial perpetuum mobiles may disappear suddenly.
IV The Kelly criterion in uncertain markets
The one-dimensional Kelly criterion is in this section extended to uncertain winning probabilities. The Kullback-Leibler (KL) distance or some more general Bregman divergence between the assumed and real probabilities captures the idea of lack of knowledge. The application of the KL distance to the Kelly criterion was already discussed in meister2024 but in the earlier paper the emphasis was on large deviations and the ability to distinguish strategies associated with different investment fractions.
Before a toy model for uncertainty and the influence on the investment fraction is discussed, we study the deliberate over- or under-investment in a situation with a known winning probability. It was shown in meister2011 how this affects the portfolio growth rate negatively as well as asymmetrically. This result can be understood intuitively, since if one underinvests, the growth underperforms but is bounded by zero. If one instead over-invests, the growth is not bounded by zero and instead can turn negative. Next, the intuition is confirmed by expanding the growth rate in terms of a small parameter linked to the optimal fraction. If the winning probability is known to be $p$ in the standard double-or-nothing game, and one varies the optimal investment fraction by an infinitesimal amount $\epsilon$, i.e. $f=2p-1+\epsilon$, then this leads to
$p\log(1+f)+(1-p)\log(1-f)=p\log(2p+\epsilon)+(1-p)\log(2-2p-\epsilon)$
$=$ $H(p)+1+p\log\Big{(}1+\frac{\epsilon}{2p}\Big{)}+(1-p)\log\Big{(}1-\frac{\epsilon}{2-2p}\Big{)}$
$=$ $H(p)+1+p\Big{(}\frac{\epsilon}{2p}-\frac{\epsilon^{2}}{8p^{2}}+\frac{\epsilon^{3}}{24p^{3}}\Big{)}-(1-p)\Big{(}\frac{\epsilon}{2-2p}+\frac{\epsilon^{2}}{8(1-p)^{2}}+\frac{\epsilon^{3}}{24(1-p)^{3}}\Big{)}+O\big{(}\epsilon^{4}\big{)}$
$=$ $H(p)+1-\frac{1}{8(1-p)p}\epsilon^{2}-\frac{2p-1}{24(1-p)^{2}p^{2}}\epsilon^{3}+O\big{(}\epsilon^{4}\big{)}.$
Let us phrase the above result in terms of the existence and sign of the quadratic and cubic terms in $\epsilon$. The growth is suboptimal for non-zero $\epsilon$, i.e. the leading term is quadratic and has a negative sign, and the loss in the portfolio growth rate due to over-investment is larger than due to under-investment for the same absolute value in $\epsilon$, i.e. the cubic term has a negative sign.
The inputs of the toy model for investment uncertainty discussed next are the value of the KL divergence $D_{\mathrm{KL}}\left(q\parallel p\right)$ set equal to $\alpha$, which is assumed to be small, i.e. $\alpha\ll 1$, as well as the probability $q$, which is assumed to be larger than $1/2$. The separation of $q$ from the unknown ‘correct’ probability is given by the KL divergence $\alpha$. To avoid possible solutions for $p$ from falling below $1/2$, which would cause the associated investment fraction to be negative, $q$ is set sufficiently above $1/2$. We further define $p=q+\epsilon$, where $\epsilon$ is small. An expansion of $D_{\mathrm{KL}}\left(q\parallel p\right)$ in terms of $\epsilon$ has the form
$D_{\mathrm{KL}}\left(q\parallel p\right)=p\log\Big{(}\frac{p}{q}\Big{)}+(1-p)\log\Big{(}\frac{1-p}{1-q}\Big{)}$
$=$ $\sum_{n=1}^{\infty}\Big{(}q+\epsilon\Big{)}(-1)^{n}\frac{\epsilon^{n}}{n\,\,q^{n}}-\Big{(}1-q-\epsilon\Big{)}\frac{\epsilon^{n}}{n\,\,(1-q)^{n}}$
$=$ $\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n}}{q^{n-1}}-\frac{1}{(1-q)^{n-1}}\Big{)}=\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n}(1-q)^{n-1}-q^{n-1}}{q^{n-1}(1-q)^{n-1}}\Big{)}$
$=$ $\frac{1}{2(1-q)q}\epsilon^{2}-\frac{1}{6}\frac{(1-q)^{2}+q^{2}}{(1-q)^{2}q^{2}}\epsilon^{3}+\frac{1}{12}\frac{(1-q)^{3}-q^{3}}{(1-q)^{3}q^{3}}\epsilon^{4}-\frac{1}{20}\frac{(1-q)^{4}+q^{4}}{(1-q)^{4}q^{4}}\epsilon^{5}+O(\epsilon^{6}).$
If, as stated above, both $D_{\mathrm{KL}}\left(q\parallel p\right)$, equal to $\alpha$, and $q$ are fixed, then in lowest approximation, as one solves the equation up to a quadratic terms, one gets two solutions for $\epsilon$. Various cases can be considered. One can consider the worst case, i.e. the minimal solution for $p$, the best case, i.e. the maximal solution for $p$, and a ‘least informative’ case, i.e. with equal weights for both possible solutions of $p$. In the equal weighted case, due to the asymmetry associated with over- and under-investment, as described above, the optimal investment fraction is lower than $12q-$ 1. To summarise, uncertainty about the right value for the winning probability leads to a lowering of the growth rate and the optimal fraction. An attractive formula for the adjusted investment fraction that captures this heuristically is $f^{*}\exp(-\lambda\alpha)$, where $f^{*}$ is the optimal uncertainty-free fraction $12q-$ 1, and $\lambda$ is a risk-aversion parameter. This result can be extended from KL divergence to the more general Bregman divergence by doing the relevant Taylor expansions for cases of interest like the Itakura–Saito distance. A similar ‘conservatism’ can be observed.
The conclusion attempts again to shed light on treasury companies.
V Conclusion
Crypto treasury companies with a rising ‘Bitcoin per share’ and mNAV above one seem to have discovered a financial perpetuum mobile and are eager to recruit followers, since copy-cats compete for the restricted supply of Bitcoins and in the process push up prices. Is there a fault in their logic? First, credit risk aggregates in the system, when these stocks are used as collateral to obtain leverage to race ahead in the accumulation process. This risk has to be managed to mitigate potential losses, when there is a dramatic price reversal of BTC or mNAV sufficiently compresses leading to portfolio liquidations. Second, if price impact of buying limited supply tokens is central to the scheme, then there is the possibility of a widening gap between the price and the value of these tokens. As Fischer Black once stated for equities, the gap between price and value in an ‘efficient market’ [^4] Fischer Black in “Noise”, Journal of Finance, vol. 41, pp. 529–543 (1986) stated that he thought ‘markets are efficient almost all of the time. “Almost all” means $90\%$ of the time’. This statement was and to a lesser extent still is controversial. is ‘almost all’ of the time not larger than $100\%$. Constraints on value exists even for crypto tokens, since a market capitalisation much beyond the current single digit trillion dollar amount has a good chance to unbalance not just the financial system and the economy in general, but also politics.
Can the gap between market capitalisation and net asset value hold indefinitely as more equity is issued, bought by current holders, and used as collateral for loans to buy further tokens? This seems unlikely, since the strategy has a weak point: Credit providers. The willingness of credit providers to lend against stocks is surely limited, since their risk appetite is bounded. What can credit providers do? They should be cognizant of the gap between market capitalisation and NAV and align their lending to the limited liquidation value under market stress of large crypto-token portfolios. It seems brokers are aware of the risk and are restricting lending against concentrated positions in token treasury companies [^5] This is suggested by Reddit discussion forums and similar anecdotal evidence.. This could break the cycle, but, as the saying goes, ‘the market can stay irrational longer than investors can stay solvent’.
The history of finance is full of wondrous schemes. After the South Sea bubble even Newton was supposed to have said, ‘I can calculate the motion of heavenly bodies, but not the madness of people’, and similar events in France, where the ‘Mississippi Bubble’ ended badly for the buccaneer John Law, undermined the established order, and prepared the ground for the coming revolution.
In the second part of the paper, simple information theoretic ideas are employed to understand the influence of uncertainty on decision making, and an asymmetric effect on the optimal fraction for the binary-choice Kelly criterion was observed as the winning probability becomes uncertain.
The paradox of high mNAV crypto treasury companies is captured by the White Queen in Lewis Carroll’s ‘Through the Looking Glass’, when she says, ‘[t]he rule is, jam to-morrow and jam yesterday – but never jam to-day’.
  1. (1) J. Kelly, A new interpretation of information rate. Bell System Technical Journal 35(4): 917–926 (1956).
  2. (2) Y. Lv & B. K. Meister, “Applications of the Kelly Criterion to multi-dimensional Diffusion Processes” International Journal of Theoretical and Applied Finance 13, 93-112. (2010) 27, reprinted in ”The Kelly Criterion: Theory & Practice:”, ed. by MacClean, Thorpe & Ziemba, World Scientific. 285-300. (2011).
  3. (3) B. K. Meister, “Meta-CTA Trading Strategies based on the Kelly Criterion”, arXiv:1610.10029. (2016).
  4. (4) B. K. Meister, “Meta-CTA Trading Strategies and Rational Market Failures”, arXiv:2209.05360. (2022).
  5. (5) B. K. Meister, “Gambling the World Away: Myopic Investors”, arXiv:2302.13994. (2023).
  6. (6) B. K. Meister, “Application of the Kelly Criterion to Prediction Markets ”, arXiv:2412.14144. (2024).
  7. (7) S. Nakamoto, ‘Bitcoin: A Peer-to-Peer Electronic Cash System’ (2008).